# Optimization Proxies using Limited Labeled Data and Training Time – A Semi-Supervised Bayesian Neural Network Approach

### v1

@@ Appendix D Concentration Bounds
Theorem 1 (Hoeffding’s).
Hoeffding (1994) Let $X_{1},\ldots,X_{M}$ i.i.d. random variables and suppose that $|X_{i}|\leqslant R$ with expectation $\mathbb{E}(X_{i})$, and let $\bar{X}_{M}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$. Then, with probability at least $1-\delta$,
@@
Theorem 2 (Empirical Bernstein).
Audibert et al. (2007); Mnih et al. (2008) Let $X_{1},\ldots,X_{M}$ be i.i.d. and suppose that $|X_{i}|\leqslant R$, expectation $\mathbb{E}(X_{i})$ and let $\bar{X}_{M}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$. With probability at least $1-\delta$,
@@
Theorem 3 (Bernstein).
Sridharan (2002) Let $X_{1},\ldots,X_{M}$ be i.i.d. and suppose that $|X_{i}|\leqslant R$, mean $\mathbb{E}(X_{i})$ and ${\mathbb{V}}=\mathrm{Var}(X_{i})$. With probability at least $1-\delta$,

### v3

@@ Appendix E Concentration Bounds
Theorem E.1 (Hoeffding’s).
(Hoeffding, 1994) Let $X_{1},\ldots,X_{M}$ i.i.d. random variables and suppose that $|X_{i}|\leqslant R$ with expectation $\mathbb{E}(X_{i})$, and let $\bar{X}_{M}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$. Then, with probability at least $1-\delta$,
@@
Theorem E.2 (Empirical Bernstein).
(Audibert et al., 2007; Mnih et al., 2008) Let $X_{1},\ldots,X_{M}$ be i.i.d. and suppose that $|X_{i}|\leqslant R$, expectation $\mathbb{E}(X_{i})$ and let $\bar{X}_{M}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$. With probability at least $1-\delta$,
@@
Theorem E.3 (Bernstein).
(Sridharan, 2002) Let $X_{1},\ldots,X_{M}$ be i.i.d. and suppose that $|X_{i}|\leqslant R$, mean $\mathbb{E}(X_{i})$ and ${\mathbb{V}}=\mathrm{Var}(X_{i})$. With probability at least $1-\delta$,
@@
Figure 8: Performance metrics for Sandwich BNN SvP model with 10-minute training.
Figure 9: Performance metrics for Supervised BNN SvP model with 10-minute training.
Figure 10: Performance metrics for Sandwich BNN model with 10-minute training.
Figure 11: Performance metrics for Supervised BNN model with 10-minute training.
Figure 12: Performance metrics for ‘Case118’ with fixed unsupervised samples (2048) over $T_{max}=900sec$ (15 minutes) using the Sandwich BNN SvP model.
Figure 13: Performance metrics for Supervised BNN SvP model model with 15-minute training.
Figure 14: Performance metrics for Sandwich BNN model with 15-minute training.
Figure 15: Performance metrics for Supervised BNN model with 15-minute training.